# Exploring Kolmogorov-Arnold Networks for Interpretable Time Series Classification

### v2

@@ V-C Hyperparameter impact analysis
We analyze the impact of key hyperparameters on KAN and Efficient KAN models, with each Fig. 4 5 6 illustrating results for specific hyperparameter variations as detailed in Section IV-D. Each figure includes line plots for five different learning rates, where the same color represents the results with the same learning rate for both models, with KAN depicted with solid lines and Efficient KAN with dashed lines.
@@
At the lowest learning rates, Efficient KAN achieved the highest F1 scores, with $\text{lr}=0.0001$ and $\text{lr}=0.001$ yielding nearly identical and stable results across grid sizes. For these rates, the F1 score exhibited a slight, linear decrease from $0.69$ at $G=3$ to approximately $0.66$ at $G=20$, suggesting minimal sensitivity to grid variation. The learning rate $\text{lr}=0.01$ yielded slightly lower performance than $\text{lr}=0.0001$ and $\text{lr}=0.001$, following a similarly gradual decrease across grid sizes. In contrast, higher learning rates ($\text{lr}=0.1$ and $\text{lr}=1$) resulted in significantly lower F1 scores, averaging around $0.3$ across all grid sizes, showing limited grid sensitivity but almost 50% reduced effectiveness relative to lower rates.
KAN, in contrast, exhibited an opposing learning rate trend, with performance improving as learning rates increased. All learning rates for KAN showed a similar pattern: a substantial drop in F1 score from $G=3$ to $G=10$, followed by a more gradual decline with increasing grid size. Notably, even KAN’s best configurations fell well below the best results of Efficient KAN, underscoring a clear performance gap. A deviation in KAN’s trend was seen at $\text{lr}=1$ for $G=3$, where the F1 score was unexpectedly lower than at $G=5$, unlike other rates where $G=3$ consistently achieved the highest performance.
@@
Model Configuration Precision Recall F1 score Time (seconds)
@@
TABLE III: Comparison of top-performing KAN models vs the reference MLP and Hive-Cote2 SotA.
@@
Out of all computed results across various hyperparameter configurations, we selected the best-performing KAN and Efficient KAN models. Their results are presented in Table III. Table is structured as Table I, along with additional column showing average training time in seconds.
@@
For comparison, we also include results from the MLP model discussed in Section V-A and Hive-Cote 2.0 as a state-of-the-art benchmark. Efficient KAN demonstrates comparable performance to MLP, with higher F1 scores and significantly faster training times. Although Hive-Cote 2.0 achieves the highest overall performance, Efficient KAN provides a balanced tradeoff with substantially reduced training time, making it a practical alternative for time-sensitive applications. This is further illustrated in the critical difference diagrams in Fig. 7, where Efficient KAN ranks above MLP but still below Hive-Cote 2.0, reflecting its competitive accuracy while maintaining efficiency. In contrast, KAN ranks lowest, showing a more significant deviation in performance.
@@ V-D Interpretability analysis
[…]
Figure 8: KAN model
@@
In this section, we evaluate the interpretability of the KAN model and compare it with the interpretability of MLP, following the methodology described in Section IV-F. The interpretability of the two models is demonstrated using the SmoothSubspace UCR dataset shown in Fig. 1. We compare the interpretability of the KAN model with the best MLP configuration (see Table III), which achieved an F1 score of almost 0.88 on the used dataset, which is similar to the fine-tuned KAN model utilised to demonstrate interpretability described in Section  IV-F.
Fig. 8 provides a visualization of the structure of the KAN model, where the thickness of the edges represents the importance of the connections based on the edge scores according to Section IV-F. The corresponding B-splines for these edges show which function is applied to the input features and provide information on how the input features are modified by the architecture. The x-axis of the enhanced B-spline in Fig. 8 represents the range of input features, while the y-axis indicates the corresponding outputs of the spline function applied to these inputs.
Looking at Fig. 8, it can be seen that features $x_{5}$, $x_{10}$, and $x_{12}$ emerge as the most influential for the models decision, based on the edge weights where thicker lines represent higher edge weight. Looking at Fig. 1 that input features $x_{5}$, $x_{10}$ are effective in differentiating class $y_{2}$ from classes $y_{1}$ and $y_{3}$, while feature $x_{12}$ further distinguishes between classes $y_{1}$ and $y_{3}$. Additionally, looking at class $y_{3}$ in Fig. 1, we can see that majority of the values within the dataset are whithin $[0,1]$ for for input feature $x_{12}$. Looking at the learned B-spline corresponding to this feature in Fig. 8 we can see that when $x_{12}$ is between 0 and 1, the spline increases the values up to 300, while the values outside of this range decrease to values between 0 and -200, diminishing their contribution. Similar observations can also be made for the input features $x_{5}$, $x_{10}$.
To further quantify feature contributions and also compare feature importance of KAN to MLP, we apply SHAP analysis as described in Section IV-F. Fig. 9 and 10 display the feature importance for each class of the SmoothSubspace dataset for both KAN and MLP, sorted from the most important to the least important feature for each predicted class. Each figure is divided into three sections, corresponding to the three output classes. In the plots, each feature’s contribution is represented by its SHAP value, which indicates whether it positively contributes to the prediction or negatively. Feature values are color-coded, with red indicating high values and blue low values. As it can be seen in Fig. 1, $2$ represents the highest value, while $-2$ represents the lowest value. Middle-range values between $-1$ to $1$, are shown as purple dots. Purple dots correspond to the flat parts of the series, where consecutive values show minimal variation, while red and blue dots are associated with the highly fluctuating sections of the series.
@@
Looking at Fig. 10, the SHAP feature importance scores are presented for the MLP. As it can be seen, feature $x_{7}$ is the most important in distinguishing classes $y_{1}$ and $y_{2}$, compared to $x_{10}$ for KAN model. However, for both MLP and KAN the $x_{10}$ and $x_{7}$ are shown as top 4 contributing features in distinguishing between the classes, indicating that both MLP and KAN model in general consider similar features as key to distinguish between the three classes.
@@ VI Conclusion
The aim of this study was to conduct a comprehensive and robust exploration of the KAN architecture for time series classification. To achieve that, we first investigated the transferability of reference Kolmogorov-Arnold Networks (KANs) from regression to classification tasks followed by key hyperparameter search on two KAN implementations in view of finding the best architecture for the task at hand and concluding with an interpretability analysis. Our findings show that the EfficientKAN implementation has a significant performance improvement over the original KAN and is also superior to MLPs by 6 percentage points in F1 score on the reference architectures. The hyperparameter analysis revealed that Efficient KAN consistently outperformed KAN across variations in grid size, depth, and layer size, showcasing its robustness and adaptability to diverse datasets. Efficient KAN proved more stable than KAN across grid sizes, depths, and layer configurations, particularly with lower learning rates. The interpretability of KANs was confirmed through SHAP analysis and B-spline visualization, highlighting their potential for transparent decision-making. In contrast, MLPs, though competitive in performance, exhibited less interpretability. Future work could explore hybrid architectures and further optimization to enhance KAN’s application scope.
Acknowledgments

### v3

@@ V-C Hyperparameter impact analysis
In this section, we analyze the impact of key hyperparameters on KAN and Efficient KAN models, with Figs. 4, 5 and 6 illustrating results for specific hyperparameter variations as detailed in Section IV-D. Each figure includes line plots for five different learning rates, where the same color represents the results with the same learning rate for both models, with KAN depicted with solid lines and Efficient KAN with dashed lines.
@@
At the lowest learning rates, Efficient KAN achieved the highest F1 scores, with $lr=0.0001$ and $lr=0.001$ yielding nearly identical and stable results across grid sizes. For these rates, the F1 score exhibited a slight, linear decrease from $0.69$ at $G=3$ to approximately $0.66$ at $G=20$, suggesting minimal sensitivity to grid variation. The learning rate $lr=0.01$ yielded slightly lower performance than $lr=0.0001$ and $lr=0.001$, following a similarly gradual decrease across grid sizes. In contrast, higher learning rates ($lr=0.1$ and lr=1\ lr=1) resulted in significantly lower F1 scores, averaging around $0.3$ across all grid sizes, showing limited grid sensitivity but almost 50% reduced effectiveness relative to lower rates.
KAN, in contrast, exhibited an opposing learning rate trend, with performance improving as learning rates increased. All learning rates for KAN showed a similar pattern: a substantial drop in F1 score from $G=3$ to $G=10$, followed by a more gradual decline with increasing grid size. Notably, even KAN’s best configurations fell well below the best results of Efficient KAN, underscoring a clear performance gap. A deviation in KAN’s trend was seen at $lr=1$ for $G=3$, where the F1 score was unexpectedly lower than at $G=5$, unlike other rates where $G=3$ consistently achieved the highest performance.
@@
Model Configuration Precision Recall F1 score Training time (seconds)
@@
TABLE V: Comparison of top-performing KAN models vs the reference MLP, InceptionTime and Hive-Cote2 SotA on 117 datasets in UCR archive.
@@
Out of all computed results across various hyperparameter configurations, we selected the best-performing KAN and Efficient KAN models. Their results are presented in Table V. The table is structured in a similar manner as Table III, along with additional column showing mean and standard deviation of average training time in seconds.
@@
For comparison, we also include results from the MLP model discussed in Section V-A, InceptionTime [32] baseline and Hive-Cote 2.0 as a state-of-the-art benchmark. Efficient KAN demonstrates comparable performance to MLP, with higher F1 scores and significantly faster training times. Although Hive-Cote 2.0 achieves the highest overall performance, Efficient KAN provides a balanced tradeoff with substantially reduced training time, making it a practical alternative for time-sensitive applications. This is further illustrated in the critical difference diagrams in Fig. 7, where Efficient KAN ranks above MLP but still below Hive-Cote 2.0 and InceptionTime, reflecting its competitive accuracy while maintaining efficiency. In contrast, KAN ranks lowest, showing a more significant deviation in performance.
(a) Efficient KAN vs KAN
(b) Efficient KAN vs MLP
(c) KAN vs MLP
Figure 8: Scatter plots comparing mean F1 scores of different model pairs.
To provide a more detailed insight into comparison between the three best developed models, we also include three scatter-plots in Fig. 8 , that visualize pairwise comparisons between them: KAN vs. MLP, Efficient KAN vs. MLP, and KAN vs. Efficient KAN. In each scatterplot, the mean F1-score across 5 different runs for each dataset obtained by one model is plotted on the x-axis, while the mean F1-score of the other model is plotted on the y-axis. These plots provide an intuitive visual comparison of relative performance across datasets. Points above the diagonal line indicate datasets where the model on the y-axis outperforms the one on the x-axis.
In Figure 8(a) , Efficient KAN is compared to the original KAN. The majority of points lie above the diagonal, indicating that Efficient KAN achieves higher F1-scores than the original KAN on most datasets. Similarly, scatter plot in Figure 8(b) confirms that Efficient KAN generally outperforms MLP. In Figure 8(c) , original KAN is compared to MLP. While MLP outperforms KAN on a number of datasets, the distinction is less prominent compared to the previous two comparisons.
@@ V-D Interpretability analysis
In this section, we evaluate the interpretability of the KAN model and compare it with the interpretability of MLP, following the methodology described in Section IV-F. The interpretability of the two models is demonstrated using the SmoothSubspace dataset from the UCR archive shown in Fig. 1. We compare the interpretability of the KAN model with the best MLP configuration (see Table V), which achieved an F1 score of almost 0.88 on the used dataset, which is similar to the fine-tuned KAN model utilised to demonstrate interpretability described in Section  IV-F.
Unlike traditional neural networks such as MLPs, which require post-hoc techniques like SHAP to gain interpretability, KANs offer interpretability by design through their composition graph. Fig. 9 provides a visualization of the structure of the KAN model, where the thickness of the edges represents the importance of the connections based on the edge scores according to Section IV-F. The corresponding B-splines for these edges show which function is applied to the input features and provide information on how the input features are modified by the architecture. The x-axis of the enhanced B-spline in Fig. 9 represents the range of input features, while the y-axis indicates the corresponding outputs of the spline function applied to these inputs.
Figure 9: KAN model with edge thickness indicating connection importance and B-spline plots showing learned functions on edges.
Looking at the KAN graph in Fig. 9, it can be seen that features $x_{5}$, $x_{10}$, and $x_{12}$ emerge as the most influential for the models decision, based on the edge weights where thicker lines represent higher edge weight. Looking at Fig. 1 that input features $x_{5}$, $x_{10}$ are effective in differentiating class $y_{2}$ from classes $y_{1}$ and $y_{3}$, while feature $x_{12}$ further distinguishes between classes $y_{1}$ and $y_{3}$. Additionally, looking at class $y_{3}$ in Fig. 1, we can see that majority of the values within the dataset are whithin $[0,1]$ for input feature $x_{12}$. Looking at the learned B-spline corresponding to this feature in Fig. 9 we can see that when $x_{12}$ is between 0 and 1, the spline increases the values up to 300, while the values outside of this range decrease to values between 0 and -200, diminishing their contribution. Similar observations can also be made for the input features $x_{5}$ and $x_{10}$. In contrast, looking at feature $x_{9}$, we can see that it provides limited relevance for classification indicated by its thinner connection edges in Fig. 9 . As seen in Fig. 1, it helps separating class $y_{2}$, but offers little distinction between classes $y_{1}$ and $y_{3}$, where values largely overlap. This aligns with the low model sensitivity reflected in the KAN graph.
@@
To validate and further quantify feature contributions and also compare feature importance of KAN to MLP, we apply SHAP analysis as described in Section IV-F. This allows us to validate whether the influential features identified through KAN’s structure and learned B-splines in Fig. 9 are also reflected consistently by the more established SHAP explanations. Figs. 10 and 11 display the feature importance for each class of the SmoothSubspace dataset for both KAN and MLP, sorted from the most important to the least important feature for each predicted class. Each figure is divided into three sections, corresponding to the three output classes. In the plots, each feature’s contribution is represented by its SHAP value, which indicates whether it positively contributes to the prediction or negatively. Feature values are color-coded, with red indicating high values and blue low values. As it can be seen in Fig. 1, $2$ represents the highest value, while 2-2 represents the lowest value. Middle-range values between 1-1 to $1$, are shown as purple dots. Purple dots correspond to the flat parts of the series, where consecutive values show minimal variation, while red and blue dots are associated with the highly fluctuating sections of the series.
@@
For the MLP model, which lacks inherent interpretability, SHAP is essential. Fig. 11 presents SHAP-based feature importances for MLP. As it can be seen, feature $x_{7}$ is the most important in distinguishing classes $y_{1}$ and $y_{2}$, compared to $x_{10}$ for KAN model. However, for both MLP and KAN the $x_{10}$ and $x_{7}$ are shown as top 4 contributing features in distinguishing between the classes, indicating that both MLP and KAN model in general consider similar features as key to distinguish between the three classes.
However, while both models highlight similar features overall, it is evident that for KAN the SHAP color coding aligns consistently with the structure of the time series, whereas in the MLP the color mapping appears more scattered and less coherent with the underlying sequential patterns, which makes its feature contributions harder to interpret in relation to the original series.
In summary, KAN provides interpretability inherently through its graph structure and learned splines. SHAP is employed primarily to validate and complement this built-in transparency. In contrast, MLP requires SHAP or similar tools to expose any interpretability, underscoring the advantage of KANs in applications where understanding model decisions is critical.
@@ VI Conclusion
The aim of this study was to conduct a comprehensive and robust exploration of the KAN architecture for time series classification. To achieve that, we first investigated the transferability of reference Kolmogorov-Arnold Networks (KANs) from regression to classification tasks , running a large number of models across 117 diverse datasets from the UCR Benchmark. We conducted hyperparameter search on two KAN implementations in view of finding the best architecture for the task at hand and concluded with an interpretability analysis. Our findings show that the Efficient KAN implementation has a significant performance improvement over the original KAN and is also superior to MLPs by 6 percentage points in F1 score on the reference architectures. The hyperparameter analysis revealed that Efficient KAN consistently outperforms KAN across variations in grid size, depth, and layer size, showcasing its robustness and adaptability to diverse datasets. Efficient KAN proved more stable than KAN across grid sizes, depths, and layer configurations, particularly with lower learning rates. The interpretability of KANs was confirmed through SHAP analysis and B-spline visualization, highlighting their potential for transparent decision-making. In contrast, MLPs, though competitive in performance, exhibited less interpretability. Future work could explore hybrid architectures and further optimization to enhance KAN’s application scope.
Funding
@@
Conflict of Interest: The authors declare that they have no conflict of interest.